OpenStudy (anonymous):

The number n is a 2-digit number. When n is divided by 10, the remainder is 9, and when n is divided by 9, the remainder is 8. What is the value of n?

hartnn (hartnn):

hINT : \(dividend = divisor * quotient+remainder.\)

OpenStudy (anonymous):

PLEASE dont give the written answer PEOPLE!

OpenStudy (sirm3d):

do you want a clue? @Areesha.1D

OpenStudy (anonymous):

yeah @hartnn i think provided it? one minute, let me try.

OpenStudy (anonymous):

I did it like this, n/10=x+9 and n/9=x+8 and then found n in both and solve for X. Where did i go wrong?

OpenStudy (anonymous):

well another way of solving is using the modulus. Are you familiar with it? let n=10a +b, \(10a+b \equiv 9 (mod 10)\) \(10a+b \equiv 8 (mod 9)\) note that \(a \equiv b (mod n)\) then \(a-b \equiv 0 (mod n)\)

OpenStudy (sirm3d):

if you divide 23 or 43 by 10, the remainder is ___ ?

hartnn (hartnn):

u were incorrect in taking both x's.....

OpenStudy (anonymous):

@Shadowys yeah i don't....

OpenStudy (anonymous):

@hartnn i know, but i was trying to find a way... trial and error ya know?

OpenStudy (sirm3d):

let's narrow down the choices. if you divide 23 or 43 by 10, the remainder is 3. So, what kind of number leaves a remainder of 9 when diided by 10?

OpenStudy (sirm3d):

in the language of fractions, \[\large \frac{ 2\color{red}3 }{ 10 }=2\frac{ \color{red}3 }{ 10 }\\ \large \frac{ 4\color{red}3 }{ 10 }=2\frac{ \color{red}3 }{ 10 }\]

OpenStudy (anonymous):

well, it's quite handy in this question. \(10a+b≡9(mod10)\) \(10a+b-9≡0(mod10)\) means that \(\frac{10a+b-9}{10} =0\) so, \(a+ \frac{b-9}{10} =0\), so since a and b are integers<10....I guess b is known... well, then do the same thing with mod 9.

OpenStudy (anonymous):

49 29 etc

OpenStudy (sirm3d):

@Shadowys solution reveals that the last digit is 9. so the choices are 19, 29, 39, up to 89

OpenStudy (anonymous):

I dont understand shadowys solution!

OpenStudy (anonymous):

I do, however, underrstand the last digit has to be 9!!

OpenStudy (sirm3d):

trial and error should solve the problem. there are a few possible choices.

OpenStudy (anonymous):

the mod form is basically a simplified dividend=... formula. \(a \equiv b (mod n)\) means a/n will have remainder b. This gets rid of solving for (x) and also can show that since remainder is b, then if we take out b from a then it can be divided perfectly. so \(a-b \equiv 0 (mod n)\) note that n has two digit, so it is made of one tens and one ones, i.e. it can be written as n=10a+b, where a,b<10, and they are positive integers. the second case is \(10a+9 \equiv 8 (mod 9)\) so, \(10a+9-8 \equiv 0 (mod 9)\) \(10a+1 \equiv 0 (mod 9)\) the only suitable value for a is....?

OpenStudy (sirm3d):

you can also say 10a + 1 = a + 1 = 0 (mod 9)

OpenStudy (anonymous):

WHAT is mod?

OpenStudy (anonymous):

mod is..well, the above expression for dividends and remainders. The explanation is above though.

OpenStudy (sirm3d):

@Shadowys solution is a foreign language =(

hartnn (hartnn):

its just he knows modular arithmetic...

OpenStudy (sirm3d):

why don't you just divide each of the numbers 19, 29, 39, up to 89 and check which leaves a remainder of 8.

hartnn (hartnn):

i wouls also use trial and error, once i know units digit is 9...just like @sirm3d said....

OpenStudy (sirm3d):

the smallest positive number that leaves a remainder of 8 when divided by 9 is 8. agree? @Areesha.1D

OpenStudy (anonymous):

lol sorry about those...kinda foreign to me too actually...just got around using it

OpenStudy (anonymous):

So trial and error it is. Thanks Guys!!!

OpenStudy (sirm3d):

@Areesha.1D if you successively add 9 to 8, what remainder will you get from each number after division by 9?

OpenStudy (anonymous):

8

OpenStudy (sirm3d):

good. so which of the numbers END in 9?

OpenStudy (anonymous):

19 29 39 etc

OpenStudy (sirm3d):

division by 9, the numbers are : 8, 17, 26, 35, 44, 53, 62, 71, 80, 89, 98, ... division by 10, the numbers are : 9, 19, 29, 39, 49, 59, 69, 79, 89, 99, ...

OpenStudy (sirm3d):

can you identify the common number in the list?

OpenStudy (anonymous):

right 89

OpenStudy (sirm3d):

BINGO.

OpenStudy (anonymous):